Find the inverse of 1 - 2x} in Q[x]/(23 - 2) using Euclids Algorithm.
Answer #1

Here basically we will take help of Euclid's algorithm and some basic rules of quotient ring.

Given quotient ring on]/(m32) [n]/603) Thus, qx++ba+c+ (x2) a, b, c E Q} Now we want to finid the inverse of [1-2n].

Note that 1+ (x3 -2) is the identity element of the given quotient ring.

Let I: (22) Now [.-2n] = 1-2x + I Let an²+brtc + L E E O [nb/I is the inverse of. 1-2n+J. le (Qu+bx+c+1) (-2+3) = 141 Now (1-ran3+(a-2b) 22 + (b-2c)n + = P(n) (132) +r(n) where اه degree r(n) _3. Now -2an 3 + (a-2b) n² + (6-2c)n+C -29 (m3-2) + (a.2bcomparing the coefficiente 9-26 = 0, C-4451 b-2C - 0 1 ㅋb=2C 29 4 26 1 ACEIt4. 1 ThW b=2c = 2 (i+4d 을 = 2(144) a= 4 + IGC ISG

Thus we are done!